OpenStudy (anonymous):

Solve: (4x-5)(3x+2)=0

OpenStudy (hoblos):

4x-5 = 0 or 3x+2=0 x=5/4 x=-2/3

OpenStudy (anonymous):

first multiply out the two terms in the parrenthasees. Do you know how to do that? Use FOIL